OpenStudy (anonymous):

Identify the correct form of the following geometric sequence. 8 , 32 , 128 , 512 , ...

OpenStudy (anonymous):

f(x) = 4x + 8 f(x) = 8(4^x - 1) f(x) = 4(8^x) f(x) = 4^x

OpenStudy (midhun.madhu1987):